OpenStudy (anonymous):

right triangle a=3 and b=3 A. what is the value of c written in simplified radical form B. what is the value of c rounded to the nearest hundredth

OpenStudy (anonymous):

are we dealing with Pythagorean theorem?

OpenStudy (anonymous):

Hello @dog1234?

OpenStudy (anonymous):

I'm assuming that it is seeing now that its a triangle, sorry!!

OpenStudy (anonymous):

can you see what I am typing now okay

OpenStudy (anonymous):

it is showing me that I am in there now I can see everything you are writing and my math question

OpenStudy (anonymous):

yes my open study wouldn't load but anyways are you aware of what Pythagorean theorem?

OpenStudy (anonymous):

I am in the chat now but am not seeing you there

OpenStudy (anonymous):

what you are not really supposed to be in the chat? if you want help on this question, just type here for now k. do you know what the Pythagorean theorem is?

OpenStudy (anonymous):

a^2 + b^ =C

OpenStudy (anonymous):

yes!! so if a=3 and b=3 then what is the new equation?

OpenStudy (anonymous):

I know that he hypothese is the longest side

OpenStudy (anonymous):

hold on gtg ill be back in a lash dont freak if u see i leave this question k?

OpenStudy (anonymous):

I came up with 9^2 + 9^2 =162

OpenStudy (anonymous):

I just am not sure if I am doing this right nor how to write this in a radical equation simplified.

OpenStudy (anonymous):

im back, lol and well yes and no??? the equation would just be replacing for now k

OpenStudy (anonymous):

so a would be replaced by what

OpenStudy (anonymous):

I think a 9

OpenStudy (anonymous):

or 81

OpenStudy (anonymous):

uhm no... it said a=3 in the beginning so what would a=

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes and if it said b=3 in the begining then b=3 as well so you have a new equation!! 3^2+3^2=c^2 got it?

OpenStudy (anonymous):

Yes but is that radical form simplified

OpenStudy (anonymous):

no no you do that at the end!!

OpenStudy (anonymous):

now would I write 9^2 +9^2 =81

OpenStudy (anonymous):

i mean 162

OpenStudy (anonymous):

no... you would now begin to break it down like an equation

OpenStudy (anonymous):

like this... 3*3=what?

OpenStudy (anonymous):